A certain weak acid, HA, has a Ka value of 1.3×10−7.

Calculate the percent ionization of HA in a 0.010 M solution.
Answer #1

HA dissociates as:

HA -----> H+ + A-

1*10^-2 0 0

1*10^-2-x x x

Ka = [H+][A-]/[HA]

Ka = x*x/(c-x)

Assuming x can be ignored as compared to c

So, above expression becomes

Ka = x*x/(c)

so, x = sqrt (Ka*c)

x = sqrt ((1.3*10^-7)*1*10^-2) = 3.606*10^-5

since c is much greater than x, our assumption is correct

so, x = 3.606*10^-5 M

% dissociation = (x*100)/c

= 3.606*10^-5*100/0.01

= 0.361 %

Answer: 0.361 %

Feel free to comment below if you have any doubts